# Interplay between Majorana and Shiba states in a minimal Kitaev chain coupled to a superconductor

### v1

@@ V Interplay between YSR and MZMs
The previous transport characteristics can be understood in terms of the Local Density of States (LDOS) which provides a way to study the transition from the weak to strong coupling to the leads ($\Gamma_{s}/t\gtrsim 1\gtrsim\Gamma_{n}/t$), resolving the degree of localization of the different subgap states, cf. App. B.
@@
The LDOS at QD2, Fig. 4(a), features a transition around $\Gamma_{s}/t\sim 0.5$ in which the excited state at $\omega=2\Delta$ disappears and re-appears as a subgap state The transition also reflects on a different localization of the states in the weak and strong coupling regimes, see Figs. 4(b,e). For weak coupling to the SC ($\Gamma_{s}/t\ll 1$), the system hosts confined MZMs where the SC lead acts as a gapped lead, Fig. 4(f). Therefore, we characterize this regime as a minimal Kitaev chain, where the excited states are just broadened by the coupling to the leads, see Fig. 4(g). In contrast, in the strong coupling regime $\Gamma_{s}/t\gtrsim 1$, both the MZM and the subgap state at $\omega\lesssim\Delta$, extend into the SC lead, Fig. 4(b, c). In this limit, the SC lead + QD1 system is described by a YSR-like state, where QD1 is subsumed in the SC lead, see Fig. 4(d). The coupling between the YSR state and QD2 renormalizes the excited state, reducing its energy while the MZM remains at $\omega=0$. In this limit, the correct physical picture is no longer that of a minimal Kitaev chain weakly coupled to both SC and normal leads, but rather of a YSR coexisting with a MZM.
@@
The energy of these YSR states decreases as $\Gamma_{s}$ increases, merging with the MZMs at zero energy for $\Gamma_{s}/t\rightarrow\infty$ for a fully spin-polarized QD1. A non-zero $\Gamma_{n}$ allows for the merging to occur at a finite value of the coupling with the SC lead, making it hard to demonstrate the Majorana properties of the chain as the gap to excited states goes to zero [43], see Fig. 5(a). The dashed line shows the convergence of the YSR state to zero energy, described by Eq. (1), and the width of the state is determined by $\Gamma_{n}$. Even though the YSR and the Majorana states overlap, the zero-bias conductance quantization is preserved (not shown).
@@
The merging of the YSR and the MZM states can be understood in terms of the LDOS across the junction and the EP associated to the zero poles, i.e., non-Hermitian spectral degeneracies in the complex spectrum of the system when considering the strong coupling to the leads. This EPs are seen as bifurcations in the poles’ imaginary part where eigenvalues and eigenvectors coalesce [41]. For $\Gamma_{s}/t\geq 1$, along with the EPs associated to the MZM, new secondary EPs appear linked to the quantized conductance behaviour of the YSR excited state in this regime, cf. App. A. Moreover, we observe a new regime at $\Gamma_{s}/t\gtrsim 4$ such there is a transition of the pole defining the dispersion of the YSR state, see also Fig. S 1 in App. A. Furthermore, in this regime the MZM at QD1 is being absorbed by the SC lead, Fig. 5(b).
@@
Figure 6: Interplay between YSRs and MZMs when partial spin-polarization is considered in QD1. (a) QD2 LDOS as a function of the hybridization parameter to the SC lead going to the ultra-strong coupling for for $E_{z}/t=10$. (b) Normal conductance for $\lambda_{s}/t=1,5,10$ marked in (a). (c,d) QD2 LDOS for $E_{z}/t=50,100$. We set $\lambda_{n}/t=1$ in all panels .
@@ VI Finite spin-polarization
In real experiments, the QDs have a finite spin polarization. Despite both QDs being subjected to the same global magnetic field, a strong coupling to the SC will strongly renormalize the g-factor of QD1 [39]. To account for this effect, we consider a finite Zeeman splitting in QD1 ($E_{z}$) while keeping QD2 spin-polarized, although our findings remain qualitatively invariant for a finite spin-splitting in QD2.
Due to the partial polarization in QD1, the coupling with the SC lead is able to mix spin up and down, leading to a competition between the Zeeman splitting and the hybridization with the SC. This competition reflects in the density of states as an avoided crossing between the MZMs and the YSR states, see Fig. 6(a). After the maximum approaching, the Majorana properties are spoiled, including the quantized zero-bias conductance, see Fig. 6(b). The transition happens for $\Gamma_{s}\propto E_{Z}$. Moreover, when the YSR state is well detached from the continuum, it exhibits quantized conductance behaviour and coupling dependent width. However, this quantized regime is reached when the ZBP is completely lost, as a signature of the destructive interference in contrast to the fully spin-polarized case.
Figure 6(c,d) reflect that as the Zeeman field in QD1 increases, the crossing point is displaced to larger values of the coupling to the SC lead in such a way that for full polarization, the regime of destructive interference is never reached and the gap to the excited state reduces with $\Gamma_{s}$ [40]. Therefore, in the ultra-strong coupling regime, the degree of polarization is relevant to the physics of the problem. In this regime MZMs are present only for $\Gamma_{s}\ll E_{z}$. Thus, it is important to preserve a well-defined spin-polarization of the YSR state along with including the effects of the continuum of quasiparticles when considering Kitaev chains formed by ABS-QDs building blocks [39, 40].
@@ Appendix A Poles Structure and EPs
In the main text, we have described a hybrid setup consisting in a quantum dot (QD) based minimal Kitaev chain coupled to a normal and a superconducting (SC) lead which can be described by the boundary Green’s functions (bGF) [46, 31, 47, 48, 49] written in the $4\times 4$ (hat notation) Nambu basis $\hat{\Psi}=(\psi_{\uparrow},\psi_{\uparrow}^{\dagger},\psi_{\downarrow},\psi_{\downarrow}^{\dagger})^{T}$
$\hat{{\cal G}}_{n}=\frac{i}{t_{n}}\tau_{0}\ ,\quad\hat{{\cal G}}_{s}=-\frac{\omega\tau_{0}-\Delta_{s}\sigma_{y}\tau_{y}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega^{2}}}\ ,$ (3)
where $\tau_{\mu}$ ($\sigma_{\mu}$) are Pauli matrices acting in spin (electron/hole) space, both in the wide-band approximation $t_{n},\,t_{s}\gg\omega,\,\Delta_{s}$. The coupling terms between the different elements across the junction take the form $\Sigma_{n/s}=\lambda_{n/s}\sigma_{z}\tau_{0}$ and, the toy model for the minimal Kitaev chain [11] in this basis,
$\hat{H}_{1/2}=\varepsilon_{1/2}\sigma_{z}\tau_{0}+2E_{z,1/2}\Pi_{\downarrow}\ ,\quad\hat{T}=(t\sigma_{z}+i\Delta_{p}\sigma_{y})\Pi_{\uparrow}\ ,$ (4)
where $\Pi_{\uparrow,\downarrow}=(\tau_{0}\pm\tau_{z})/2$ are the projectors over the spin up/down sector. Notice that as one reaches the polarized limit of the minimal Kitaev chain when $E_{z,1/2}\rightarrow\infty$, the coupling between the QDs in the spin down sector goes correspondingly to zero. It should be reminded that the effective coupling between QDs is obtained from a Schrieffer–Wolff transformation and therefore, it is inversely proportional to the energy difference between states. Thus, the GFs projected over the minimal Kitaev chain can be computed working directly on a spinless $2\times 2$ basis $\Psi=(\psi_{\uparrow},\psi_{\uparrow}^{\dagger})^{T}$, equivalent to remove the anomalous part of the SC lead’s GF in the calculation. This effect is associated to the incompatibility between order parameters between the SC lead and the Kitaev chain. Unless otherwise stated, we use $\Delta_{p}=\Delta_{s}=\Delta=t$ and $t_{s}=t_{n}=10t$. In the spin-less basis it is satisfied that
${\cal G}_{n}=\frac{i}{t_{n}}\sigma_{0}\ ,\quad{\cal G}_{s}=-\frac{\omega\sigma_{0}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega^{2}}}\ ,$ (5)
@@
$G_{11}={}$ $\big{[}[{\cal G}_{11}]^{-1}-\Sigma_{s}^{\dagger}{\cal G}_{s}\Sigma_{s}\big{]}^{-1},$ (6a)
$G_{22}={}$ $\big{[}[{\cal G}_{22}]^{-1}-\Sigma_{n}{\cal G}_{n}\Sigma_{n}^{\dagger}\big{]}^{-1},$ (6b)
where the bGF associated to both ends of the minimal Kitaev are
${\cal G}_{11}={}$ $[\omega\mathbb{I}-H_{1}-T{\cal G}^{(0)}_{22}T^{\dagger}]^{-1},$ (7a)
${\cal G}_{22}={}$ $[\omega\mathbb{I}-H_{2}-T^{\dagger}{\cal G}^{(0)}_{11}T]^{-1},$ (7b)
@@
${\cal G}^{(0)}_{11}={}$ $[\omega\mathbb{I}-H_{1}-\Sigma_{s}^{\dagger}{\cal G}_{s}\Sigma_{s}]^{-1},$ (8a)
${\cal G}^{(0)}_{22}={}$ $[\omega\mathbb{I}-H_{2}-\Sigma_{n}{\cal G}_{n}\Sigma_{n}^{\dagger}]^{-1}.$ (8b)
Without any loss of generality, the poles of the system (shared by both QDs) can be obtained from $G_{11}$ in Eq. (6). Then, the associated self energy due to the coupling to QD2 including the adjacent normal lead $\Sigma=T{\cal G}^{(0)}_{22}T^{\dagger}$, simplifies to
$\Sigma=\begin{pmatrix}t^{2}\Omega_{-}+\Delta_{p}^{2}\Omega_{+}&&-t\Delta_{p}(\Omega_{+}+\Omega_{-})\\ -t\Delta_{p}(\Omega_{+}+\Omega_{-})&&t^{2}\Omega_{+}+\Delta_{p}^{2}\Omega_{-}\end{pmatrix},$ (9)
@@
Figure S 1: (a) Real and imaginary part (straight and dotted lines respectively) of the poles as a function of the hybridization parameter to the SC lead obtained from Eq. (1) when considering $\Gamma_{n}\neq 0$. (b, c) EPs for $\lambda_{s}/t=5$ ($\lambda_{s}/t=2$) marked by arrows in (a). Insets show a close-up image of the fundamental EPs. We set $\lambda_{n}/t=1$ in all panels.
@@
$G_{11}=\frac{\tau\,\gamma_{+}\gamma_{-}}{P(\omega)}\begin{pmatrix}g_{e}(\omega)&&f(\omega)\\ f(\omega)&&g_{h}(\omega)\end{pmatrix},$ (10)
@@
$f(\omega)$ $=$ $-(\gamma_{+}+\gamma_{-})t\Delta_{p}\tau,$ (11)
@@
and $\tau=\sqrt{\Delta_{s}^{2}-\omega^{2}}$, where $\Gamma_{s}=\lambda_{s}^{2}/t_{s}$ is the hybridization with the SC lead. Notice that the characteristic polynomial $P(\omega)=g_{e}g_{h}-f^{2}$ has 16th degree in $\omega$ and contains all the information concerning the poles of the system. In the sweet spot and $\Gamma_{n}\rightarrow 0$ limit, the polynomial simplifies to
$P(\omega)=\omega^{4}\big{[}\big{(}\Gamma_{s}^{2}\omega^{2}+(\Delta_{s}^{2}-\omega^{2})(\omega^{2}-4\Delta_{p}^{2})\big{)}^{2}-4\Gamma_{s}^{2}(\Delta_{s}^{2}-\omega^{2})(\omega^{2}-2\Delta_{p}^{2})^{2}\big{]}.$ (12)
Along with the zero modes, there are other subgap poles that can be computed analytically as
$z^{(\mu)}_{s,e}=$ $\,\frac{1}{\sqrt{6}}\sqrt{2(8\Delta_{p}^{2}+\Delta_{s}^{2}-\Gamma_{s}^{2})+z_{\mu}^{*}\alpha^{1/3}+z_{\mu}\frac{(4\Delta_{p}^{2}-\Delta_{s}^{2})^{2}-2(8\Delta_{p}^{2}+\Delta_{s}^{2})\Gamma_{s}^{2}+\Gamma_{s}^{4}}{\alpha^{1/3}}}.$ (13a)
$\alpha=$ $\,(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3}+3(40\Delta_{p}^{4}+4\Delta_{p}^{2}\Delta_{s}^{2}+\Delta_{s}^{4})\Gamma_{s}^{2}-3(8\Delta_{p}^{2}+\Delta_{s}^{2})\Gamma_{s}^{4}+\Gamma_{s}^{6}+12\sqrt{3}\beta^{1/3},$ (13b)
$\beta^{2}=$ $\,\Delta_{p}^{4}\Gamma_{s}^{2}\big{[}(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3}+2(-2\Delta_{p}^{4}+10\Delta_{p}^{2}\Delta_{s}^{2}+\Delta_{s}^{4})\Gamma_{s}^{2}-\Delta_{s}^{2}\Gamma_{s}^{4}\big{]},$ (13c)
where the parameter $z_{\mu}$ can be either $z_{1}=-2$ or $z_{2}=(1+i\sqrt{3})$. Depending on the parameter regime, the excited state of the system ($z_{s,e}$) may be described by one of these subgap poles.
@@
Figure S 1 shows the real and imaginary parts of the poles of the system including non-Hermitian effects such $\Gamma_{n}\neq 0$. Thus, we are able to study the Exceptional Points (EPs) [42, 43, 44, 45] for different values of the coupling to the SC lead. Figure S 1(a) shows several transitions where the real part of some poles goes to zero. First, a transition at $\Gamma_{s}/t=1$ appears as one of the poles detaching from the continuum $E=\Delta$ goes to zero (i.e., $z=\sqrt{t^{2}-\Gamma_{s}^{2}}\rightarrow 0$), defining the change of regime from weak to strong coupling.
@@
In the weak coupling limit, there are only EPs associated to the Majorana Zero Modes (MZMs), where the excited state can be found at $z_{s,e}=2\Delta$. Figure S 1(c) illustrates the EPs associated to the imaginary part of the zero poles in this regime. The maximum (minimum) value of the bifurcation at the sweet spot is $\Gamma^{+}=\Gamma_{n}$ ($\Gamma^{-}=0$) signaling that one of the Majorana poles is non-decaying and completely decoupled from the leads [42, 44] (i.e., there is no effective coupling with the SC lead in the weak coupling owing to the SC gap).
@@
Above the transition, the excited state behaves as a Yu-Shiba-Rusinov (YSR) state defined by the subgap pole $z_{s,e}=z^{(1)}_{s,e}$ in Eq. (13a). Remarkably, in this regime appears other secondary EPs coexisting with the main one, see Fig. S 1(b). These secondary EPs can be related to the quantized transport behaviour of the YSR in the normal conductance mentioned in the main text. The maximum (minimum) value of the bifurcation at the sweet spot is $\Gamma^{+}=\Gamma_{s}$ ($\Gamma^{-}=\{\Gamma_{n},0\}$), signalling a configuration with maximum coupling asymmetry [42].
Figure S 2: (a) QD2 LDOS as a function of the hybridization parameter to the SC lead going to the ultra-strong coupling for $\varepsilon_{2}/t=0.5$. (b-c) LDOS as a function of QD2 level for $\lambda_{s}/t=10,40$ marked by arrows in (a). (d-e) LDOS across the junction, showing the spreading of the states into the SC lead. White dashed lines represents the pole obtained computationally from the characteristic polynomial in Eq. (10). We set $\lambda_{n}/t=1$ in all panels.
@@
Finally, Fig. S 1(a) shows another transition when $\operatorname{\mathfrak{Re}}z^{(1)}_{s,e}\rightarrow 0$ around $\Gamma_{s}/t\sim 4$, where a bifurcation of the imaginary parts of the poles take place defining the ultra-strong coupling regime. Above this transition, the excited state is described by $z_{s,e}=z^{(2)}_{s,e}$ in Eq. (13a). Finally, in the limit $\Gamma_{s}\rightarrow\infty$ the YSR state merges with the MZMs. However, when detuning QD2 the ultra-strong coupling regime physics are strongly renormalized.

### v3

@@ V Interplay between YSR and MZMs
The previous transport characteristics can be understood in terms of the Local Density of States (LDOS) which provides a way to study the transition from the weak to strong coupling to the leads (i.e., $\Gamma_{s}/t\gtrsim 1\gtrsim\Gamma_{n}/t$), resolving the degree of localization of the different subgap states, cf. App. B for details.
@@
The LDOS at QD2 Fig. 4(a), features a transition around $\Gamma_{s}/t\sim 0.5$ in which the excited state $z_{s,e}=2\Delta$ disappears and re-appears as a subgap state. The transition also reflects on a different localization of the states in the strong and weak coupling regimes, see Figs. 4(b,e) respectively. For weak coupling to the SC ($\Gamma_{s}/t\ll 1$), the system hosts confined MZMs where the SC acts as a gapped lead, Fig. 4(f). Therefore, we characterize this regime as a minimal Kitaev chain, where the excited states are just broadened by the coupling to the leads, see Fig. 4(g). In contrast, in the strong coupling regime $\Gamma_{s}/t\gtrsim 1$, both the MZM and the subgap state at $\omega\lesssim\Delta$, extend into the SC lead, Fig. 4(c). In this limit, the SC lead + QD1 system is described by a YSR-like state, where QD1 is subsumed in the SC lead, see Fig. 4(d). The coupling between the YSR state and QD2 renormalizes the excited state, reducing its energy, while the MZM remains at $\omega=0$. Thus, the correct physical picture is no longer that of a minimal Kitaev chain weakly coupled to both SC and normal leads, but rather of a YSR coexisting with a MZM.
@@
The energy of these YSR states decreases as $\Gamma_{s}$ increases, merging with the MZMs at zero energy for $\Gamma_{s}/t\rightarrow\infty$ for a fully spin-polarized QD1. A non-zero $\Gamma_{n}$ allows for the merging to occur at a finite value of the coupling with the SC lead, making it hard to demonstrate the Majorana properties of the chain as the gap to excited states goes to zero [61], see Fig. 5(a). The dashed line shows the convergence of the YSR state to zero energy, described by Eq. (3), and the width of the state is determined by $\Gamma_{n}$. Even though the YSR and the Majorana states overlap, the oscillatory behaviour of the YSR is lost [76] and the zero-bias conductance quantization is preserved, cf. App. G. When considering $\varepsilon_{1}\neq 0$ or $t/\Delta\neq 0$, the subgap states also tend towards zero energy in the ultra-strong coupling regime, however a conductance dip at zero-bias is observed associated to a destructive interference, cf. App. G.
@@
The merging of the YSR and the MZM states can be understood in terms of the LDOS across the junction and the EP associated to the zero poles, i.e., non-Hermitian spectral degeneracies in the complex spectrum of the system when considering the strong coupling to the leads. These EPs are seen as bifurcations in the poles’ imaginary part, where eigenvalues and eigenvectors coalesce [57]. For $\Gamma_{s}/t\geq 1$, along with the EP associated to the MZM, a new secondary EP appears linked to the quantized conductance behaviour of the YSR excited state in this regime, cf. App. A. Moreover, we observe a new regime at $\Gamma_{s}/t\gtrsim 4$ such there is a transition of the pole defining the dispersion of the YSR state, see also Fig. S 1 in App. A. Furthermore, in this regime the MZM at QD1 is being absorbed by the SC lead, Fig. 5(b).
@@
Figure 6: Interplay between YSRs and MZMs when partial spin-polarization is considered in QD1. (a) QD2 LDOS as a function of the hybridization parameter to the SC lead going to the ultra-strong coupling for $E_{z}/t=10$. (b) Normal conductance for $\lambda_{s}/t=1,5,10$ marked in (a). (c,d) QD2 LDOS for $E_{z}/t=50,100$. White dashed line represents the computational pole obtained from the characteristic polynomial when considering $\Gamma_{n},E_{z}\neq 0$. We set $\lambda_{n}/t=1$ in all panels.
@@ VI Finite spin-polarization
In real experiments, the QDs have a finite spin polarization. Despite both QDs being subjected to the same global magnetic field, a strong coupling to the SC will strongly renormalize the g-factor of QD1 [52]. To account for this effect, we consider a finite Zeeman splitting in QD1 while keeping QD2 spin-polarized (i.e., $E_{z,1}=E_{z}$, and $E_{z,2}\rightarrow\infty$), although our findings remain qualitatively invariant for a finite spin-splitting in QD2.
Due to the partial polarization in QD1, the coupling with the SC lead is able to mix spin up and down, leading to a competition between the Zeeman splitting and the hybridization with the SC. This competition reflects in the density of states as an avoided crossing between the MZMs and the YSR states, see Fig. 6(a). After the maximum approaching, the Majorana properties are spoiled, including the quantized zero-bias conductance, see Fig. 6(b). The transition happens for $\Gamma_{s}\sim E_{Z}$. Moreover, when the YSR state is well detached from the continuum, it exhibits quantized conductance behaviour and coupling dependent width. However, this quantized regime is reached when the ZBP is completely lost, as a signature of the destructive interference in contrast to the fully spin-polarized case, cf. App. C for details on the topological analysis in this regime.
Figure 6(c,d) reflect that as the Zeeman field in QD1 increases, the crossing point is displaced to larger values of the coupling to the SC lead in such a way that for full polarization, the regime of destructive interference is never reached and the gap to the excited state reduces with $\Gamma_{s}$ [53]. Therefore, in the ultra-strong coupling regime, the degree of polarization is relevant to the physics of the problem. In this regime, MZMs are present only for $\Gamma_{s}\ll E_{z}$. Thus, it is important to preserve a well-defined spin-polarization of the YSR state along with including the effects of the continuum of quasiparticles when considering Kitaev chains formed by ABS-QDs building blocks [52, 53].
@@ Appendix A Poles Structure and EPs
In the main text, we have described a hybrid setup consisting in a quantum dot (QD) based minimal Kitaev chain coupled to a normal and a superconducting (SC) lead which can be described by the advanced boundary Green’s functions (bGF) [64, 34, 65, 66, 67] written in the $4\times 4$ (hat notation) Nambu basis $\hat{\Psi}=(\psi_{\uparrow},\psi_{\uparrow}^{\dagger},\psi_{\downarrow},\psi_{\downarrow}^{\dagger})^{T}$,
$\hat{{\cal G}}^{A}_{n}=\frac{i}{t_{n}}\sigma_{0}\tau_{0}\ ,\quad\hat{{\cal G}}^{A}_{s}=-\frac{\omega\sigma_{0}\tau_{0}-\Delta_{s}\sigma_{y}\tau_{y}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega^{2}}}\ ,$ (S 1)
where $\tau_{\mu}$ ($\sigma_{\mu}$) are Pauli matrices acting in spin (electron/hole) space, both in the wide-band approximation $t_{n},\,t_{s}\gg\omega,\,\Delta_{s}$. For the advanced GFs we are implicitly assuming $\omega\rightarrow\omega-i\eta$ with $\eta\rightarrow 0^{+}$. Following Ref. [34] we define the brunch cut in the square root along the negative axis,
$\sqrt{\Delta_{s}^{2}-(\omega-i\eta)^{2}}=\left\{\begin{array}[]{cl}\sqrt{\Delta_{s}^{2}-\omega^{2}}\,,&|\omega|\leq\Delta_{s}\,,\\ i\,\textrm{sgn}(\omega)\sqrt{\omega^{2}-\Delta_{s}^{2}}\,,&|\omega|>\Delta_{s}\,.\end{array}\right.$ (S 2)
From now on, we omit the super-index $A$.
The coupling terms between the different elements across the junction take the form $\hat{\Sigma}_{n/s}=\lambda_{n/s}\sigma_{z}\tau_{0}$ and, the toy model for the minimal Kitaev chain [12] written in this basis,
$\hat{H}_{\nu}=\varepsilon_{\nu}\sigma_{z}\tau_{0}+2E_{z,\nu}\sigma_{z}\Pi_{\downarrow}\ ,\quad\hat{T}=(t\sigma_{z}+i\Delta_{p}\sigma_{y})\Pi_{\uparrow}\ ,$ (S 3)
where $\nu={1,2}$ corresponds to the different sites in the chain, and $\Pi_{\uparrow,\downarrow}=(\tau_{0}\pm\tau_{z})/2$ are the projectors over the spin up/down sector. Notice that as one reaches the polarized limit of the minimal Kitaev chain when $E_{z,\nu}\rightarrow\infty$, the coupling between the QDs in the spin down sector goes correspondingly to zero. It should be reminded that the effective coupling between QDs is obtained from a Schrieffer–Wolff transformation and therefore, it is inversely proportional to the energy difference between states. Thus, the GFs projected over the minimal Kitaev chain can be computed working directly on a spinless $2\times 2$ basis $\Psi=(\psi_{\uparrow},\psi_{\uparrow}^{\dagger})^{T}$, equivalent to remove the anomalous part of the SC lead’s GF in the calculation. This effect is associated to the incompatibility between order parameters between the SC lead and the Kitaev chain [69, 33, 34]. Unless otherwise stated, we use $\Delta_{p}=\Delta_{s}=\Delta=t$ and $t_{s}=t_{n}=10\,t$. In the spin-less basis it is satisfied that
${\cal G}_{n}=\frac{i}{t_{n}}\sigma_{0}\ ,\quad{\cal G}_{s}=-\frac{\omega\sigma_{0}}{t_{s}\sqrt{\Delta_{s}^{2}-\omega^{2}}}\ ,$ (S 4)
@@
$G_{11}={}$ $\big{[}[{\cal G}_{11}]^{-1}-\Sigma_{s}^{\dagger}{\cal G}_{s}\Sigma_{s}\big{]}^{-1},$ (S 5a)
$G_{22}={}$ $\big{[}[{\cal G}_{22}]^{-1}-\Sigma_{n}{\cal G}_{n}\Sigma_{n}^{\dagger}\big{]}^{-1},$ (S 5b)
where the bGF associated to both ends of the minimal Kitaev - including both QD in the chain and the opposite end lead - are
${\cal G}_{11}={}$ $[\omega\mathbb{I}-H_{1}-T{\cal G}^{(0)}_{22}T^{\dagger}]^{-1},$ (S 6a)
${\cal G}_{22}={}$ $[\omega\mathbb{I}-H_{2}-T^{\dagger}{\cal G}^{(0)}_{11}T]^{-1},$ (S 6b)
@@
${\cal G}^{(0)}_{11}={}$ $[\omega\mathbb{I}-H_{1}-\Sigma_{s}^{\dagger}{\cal G}_{s}\Sigma_{s}]^{-1},$ (S 7a)
${\cal G}^{(0)}_{22}={}$ $[\omega\mathbb{I}-H_{2}-\Sigma_{n}{\cal G}_{n}\Sigma_{n}^{\dagger}]^{-1},$ (S 7b)
being the bGF of each QD in the chain coupled to its adjacent lead.
Without any loss of generality, the poles of the system (shared by both QDs) can be obtained from $G_{11}$ in Eq. (S 5). Then, the associated self energy due to the coupling to QD2 including the adjacent normal lead $\Sigma=T\,{\cal G}^{(0)}_{22}T^{\dagger}$, simplifies to
$\Sigma=\begin{pmatrix}t^{2}\Omega_{-}+\Delta_{p}^{2}\Omega_{+}&&-t\Delta_{p}(\Omega_{+}+\Omega_{-})\\ -t\Delta_{p}(\Omega_{+}+\Omega_{-})&&t^{2}\Omega_{+}+\Delta_{p}^{2}\Omega_{-}\end{pmatrix},$ (S 8)
@@
Figure S 1: (a) Real and imaginary part (solid and dotted lines respectively) of the poles as a function of the hybridization parameter to the SC lead obtained from Eq. (S 11) when considering $\Gamma_{n}\neq 0$. (b, c) EPs for $\lambda_{s}/t=5$ ($\lambda_{s}/t=2$) marked by arrows in (a). Insets show a close-up image of the fundamental EPs. We set $\lambda_{n}/t=1$ in all panels, except indicated, we have set $t/\Delta=1$.
@@
$G_{11}=\frac{\tau\,\gamma_{+}\gamma_{-}}{P(\omega)}\begin{pmatrix}g_{e}(\omega)&&f(\omega)\\ f(\omega)&&g_{h}(\omega)\end{pmatrix},$ (S 9)
@@
$f(\omega)$ $=$ $-(\gamma_{+}+\gamma_{-})t\Delta_{p}\tau\,,$ (S 10)
@@
and $\tau=\sqrt{\Delta_{s}^{2}-\omega^{2}}$, where $\Gamma_{s}=\lambda_{s}^{2}/t_{s}$ is the hybridization with the SC lead. Notice that the characteristic polynomial $P(\omega)=g_{e}g_{h}-f^{2}$ has 8th degree in $\omega$ and contains all the information concerning the poles of the system. In the sweet spot and $\Gamma_{n}\rightarrow 0$ limit, the polynomial simplifies to
$P(\omega)=\omega^{4}\big{(}\Gamma_{s}+\sqrt{\Delta_{s}^{2}-\omega^{2}}\big{)}\left[\omega^{2}\Gamma_{s}+(\omega^{2}-4\Delta_{p}^{2})\sqrt{\Delta_{s}^{2}-\omega^{2}}\right].$ (S 11)
Along with the trivial zeros and $z_{0}=\sqrt{\Delta_{s}^{2}-\Gamma_{s}^{2}}$, the rest of the poles can be computed analytically as,
$z_{\mu}=$ $\,\frac{1}{\sqrt{6}}\sqrt{2(8\Delta_{p}^{2}+\Delta_{s}^{2}-\Gamma_{s}^{2})+\xi_{\mu}^{*}\alpha^{1/3}(\Gamma_{s})+\xi_{\mu}\frac{(4\Delta_{p}^{2}-\Delta_{s}^{2})^{2}-2\Gamma_{s}^{2}(8\Delta_{p}^{2}+\Delta_{s}^{2})+\Gamma_{s}^{4}}{\alpha^{1/3}(\Gamma_{s})}}\;.$ (S 12a)
$\alpha(\Gamma_{s})=$ $\,(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3}+3\Gamma_{s}^{2}(40\Delta_{p}^{4}+4\Delta_{p}^{2}\Delta_{s}^{2}+\Delta_{s}^{4})-3\Gamma_{s}^{4}(8\Delta_{p}^{2}+\Delta_{s}^{2})+\Gamma_{s}^{6}-12\sqrt{3}\,\beta(\Gamma_{s})\,,$ (S 12b)
$\beta(\Gamma_{s})=$ $\,\Gamma_{s}\Delta_{p}^{2}\sqrt{(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3}-2\Gamma_{s}^{2}(2\Delta_{p}^{4}-10\Delta_{p}^{2}\Delta_{s}^{2}-\Delta_{s}^{4})-\Gamma_{s}^{4}\Delta_{s}^{2}}\,,$ (S 12c)
where the parameter $\xi_{\mu}$ can be either $\xi_{1}=-2$ or $\xi_{2}=(1+i\sqrt{3})$. The subset of non-trivial poles of the system are contained in $\{z_{0},z_{1},z_{2}\}$, along with their symmetry partners $z\rightarrow-z^{*}$. It should be noted that by the choice of advanced GFs, all the physical poles satisfy $\operatorname{\mathfrak{Im}}z\geq 0$. Depending on the parameter regime, the excited state of the system ($z_{s,e}$) may be described by one of these subgap poles. In the main text we reproduce the lowest order in the superconducting hybridization of $z_{1}$ to give a comprehensible compact approximation of the excited state, i.e., $\alpha\rightarrow(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3}-12\sqrt{3}\,\beta$, and $\beta\rightarrow\Gamma_{s}\Delta_{p}^{2}(4\Delta_{p}^{2}-\Delta_{s}^{2})^{3/2}$.
@@
Figure S 1 shows the real and imaginary parts of the poles of the system [60], including non-Hermitian effects such $\Gamma_{n}\neq 0$. Thus, we are able to study the Exceptional Points (EPs) [58, 59, 60, 61, 62, 63] for different values of the coupling to the SC lead. Figure S 1(a) shows several transitions where the real part of some poles goes to zero. First, a transition at $\Gamma_{s}/t=1$ appears as one of the poles detaching from the continuum $E=\Delta$, goes to zero (i.e., $\operatorname{\mathfrak{Re}}z_{0}\rightarrow 0$), defining the change of regime from weak to strong coupling.
@@
In the weak coupling limit, there are only EPs associated to the Majorana Zero Modes (MZMs), where the excited state can be found at $z_{s,e}=2\Delta$. Figure S 1(c) illustrates the EP associated to the imaginary part of the zero poles in this regime. The maximum (minimum) value of the bifurcation at the sweet spot is $\Gamma^{+}_{1}=\Gamma_{n}$ ($\Gamma^{-}_{1}=0$) signaling a configuration with maximum coupling asymmetry [59], i.e., one of the MZM is non-decaying and completely decoupled from the SC lead [58, 59, 62] owing to the SC gap in the weak coupling.
@@
Above the transition, the excited state behaves as a Yu-Shiba-Rusinov (YSR) state defined by the subgap pole $z_{s,e}=z_{1}$ in Eq. (S 12a). Remarkably, in this regime appears other secondary EP coexisting with the main one, see Fig. S 1(b). This secondary EP can be related to the quantized transport behaviour of the YSR in the normal conductance mentioned in the main text. The maximum (minimum) value of the bifurcation at the sweet spot for the secondary EP is $\Gamma^{+}_{2}\approxeq\Gamma_{s}$ ($\Gamma^{-}_{2}=\Gamma_{n}$).
@@
Finally, Fig. S 1(a) shows another transition when $\operatorname{\mathfrak{Re}}z_{1}\rightarrow 0$ around $\Gamma_{s}/t\sim 4$, where a bifurcation of the imaginary parts of the poles take place, defining the ultra-strong coupling regime. Above this transition, the excited state is described by $z_{s,e}=z_{2}$ in Eq. (S 12a). Finally, in the limit $\Gamma_{s}\rightarrow\infty$, the YSR state merges with the MZMs.